ricks age plus sherrys age is 52 years. ricks age is double sherryys age plus 10 years. how old is each person
Answer by ankor@dixie-net.com(22740)

let r = Rick's age
let s =Sherry's
:
Write an equation for each statement
:
ricks age plus sherrys age is 52 years.
r + s = 52
ricks age is double sherry's age plus 10 years.
r = 2s + 10
in the first equation, replace r with (2s+10)
(2s + 10) + s = 52
3s = 52 - 10
3s = 42
s = 42/3
s = 14 yrs is Sherry's age
:
You can find Rick's age, replace s with 14 in either equation
:
how old is each person